The radioisotope ¹⁴C (half life = 5730 years) is used for carbon dating. What is the first order rate constant for ¹⁴C?
Answer #1

all radioactive decay are first order kinetics

for first order

t1/2 = 0.693 /K

K = 0.693 / t1/2

t1/2 = 5730 yeas

K = 0.693 / 5730

K = 1.2 x 10-4 year-1